OpenStudy (anonymous):

Does this Equation have a identity or a Not a Solution? 2(a-4)=4a-(2a+4)

OpenStudy (precal):

2a-8=4a-2a-4 2a-8=2a-4 -8=-4 False therefore no solution

OpenStudy (anonymous):

Thanks Medal Awarded ! Woot Woot!